# Forecasting age distribution of deaths across countries: Life expectancy and annuity valuation

### v2

@@ 4.1 Expanding window
We implement an expanding-window scheme, which is commonly used to evaluate model and parameter stability, as well as prediction accuracy over time. The expanding window analysis determines the stability of a model’s parameter estimates and their forecasts over an expanding window through the sample. Since each data set has a different sample period, we keep the last 20 years of data for evaluation and the remaining data for training. For the 20 years of data in the testing sample, we consider one-step-ahead to 20-step-ahead forecasts to assess the model’s performance at the short to medium horizons. Through the expanding-window scheme, we produce 20 one-step-ahead forecasts, 19 two-step-ahead forecasts, $\dots$, and one 20-step-ahead forecast. We compare these forecasts with the holdout data in the testing sample to determine the accuracy of the out-of-sample forecast. In Figure 3, we present a concept diagram.
@@ 4.2 Point forecast evaluation metrics
Since the age distribution of death counts can be considered a probability density function, we consider two density evaluation metrics. The metrics are discrete Kullback-Leibler divergence (KLD) (Kullback & Leibler, 1951) and Jensen-Shannon divergence (JSD) (Shannon, 1948).
The KLD measures information loss by approximating an unknown density with its approximation. For two probability density functions, denoted by $d_{n+\xi}(u)$ and $\widehat{d}_{n+\xi|n}(u)$, the discrete KLD is defined as
@@
$=\$$\frac{1}{111\times(21-h)}\sum^{20}_{\xi=h}\sum^{111}_{x=1}d_{n+\xi,x}\cdot(\ln d_{n+\xi,x}-\ln\widehat{d}_{n+\xi|n,x})+$
$\frac{1}{111\times(21-h)}\sum^{20}_{\xi=h}\sum^{111}_{x=1}\widehat{d}_{n+\xi|n,x}\cdot(\ln\widehat{d}_{n+\xi|n,x}-\ln d_{n+\xi,x}),$
where $\xi$ denotes the forecasting period. A feature of the discrete KLD is its non-negativity.
Even though we have a sample of observed densities in the forecasting period, they may not be the actual densities. An alternative is given by the JSD, which can be seen as a symmetric and smoothed version of the KLD. The JSD is defined by
@@
where $\delta_{n+\xi,x}$ measures a common quantity between $d_{n+\xi,x}$ and $\widehat{d}_{n+\xi,x}$. One example of $\delta_{n+\xi,x}$ can be its geometric mean $\delta_{n+\xi,x}=\sqrt{d_{n+\xi,x}\widehat{d}_{n+\xi,x}}$.
@@ 4.3 Comparison of point forecast accuracy
In Figure 3, we demonstrate the forecasts of the life-table death counts obtained from the CLR and CDF transformations. For one-step-ahead forecasts, the differences between the two transformation are marginal. As forecast horizon increases to $h=20$, the difference is more visible. As measured by the KLD and JSD, the CDF transformation produces smaller forecast errors than those from the CLR transformation for $h=1$ and $h=20$ considered.
Figure 3: One-step-ahead and 20-step-ahead forecasts of life-table death counts for Australia women (first row) and men (second row).
Averaging over the point forecast errors obtained from the 24 countries, Figure 4 presents horizon-specific plots to compare the accuracy of the two transformations, as measured by the KLD and JSD. Between the CLR and CDF transformations, the CDF transformation consistently outperforms the CLR transformation. Such superiority is credited to the monotonicity constraint enjoyed by the CDF transformation. For modeling the unconstrained data in the Euclidean space, it is advantageous to use $K=6$ for forecasting. In contrast, the number of components selected by the EVR tends to be fewer than the optimal. Choosing a larger number of components results in only a minor loss, whereas selecting too few components leads to a more substantial loss.
Figure 4: Comparison of the point forecast accuracy, measured by the KLD and JSD, between the CLR and CDF transformations. For each transformation, we implement a principal component analysis where the number of components is selected by the EVR criterion or fitted at $K=6$.
@@ 4.4 Construction of pointwise prediction intervals
Using the data in the training sample, we carry out an expanding window approach to obtain the $h$-step-ahead density forecasts in the validation set for $h=1,\dots,19$. For different forecast horizons, we have different numbers of curves in the validation set. For $h=1$, we have 20 years in evaluate the forecast errors; when $h=19$, we have two years to evaluate the residual functions between the holdout data in the validation set and their forecasts. Based on these residuals, we can compute the functional standard deviation, which requires at least two years of data. The forecast errors can be denoted by $\widehat{\epsilon}_{\nu,x}=d_{\nu,x}-\widehat{d}_{\nu,x}$, where $\widehat{d}_{\nu,x}$ denotes the forecasts obtained from the two transformations, and $\nu=1,\dots,(21-h)$ denotes the number of observations in the validation set for $h=1,\dots,19$. When $h=1$, we have 20 observations in the validation set; when $h=19$, we have two observations.
Let us denote $\Gamma_{\nu,x}=\text{sd}(\widehat{\epsilon}_{\nu,x})$, where $\text{sd}(\cdot)$ can be pointwise standard deviation. We denote $\bm{\Gamma}_{\nu}=(\Gamma_{\nu,1},\dots,\Gamma_{\nu,111})^{\top}$ and $\bm{\widehat{\epsilon}}_{\nu}=(\widehat{\epsilon}_{\nu,1},\dots,\widehat{\epsilon}_{\nu,111})^{\top}$. For a level of significance $\alpha$, customarily $\alpha=0.2$ or 0.05, we aim to find $\theta_{h,\alpha}$ such that approximately $100(1-\alpha)\%$ of the residuals satisfy
@@
By the law of large numbers, one may achieve
@@
To determine the optimal $\theta_{h,\alpha}$, the samples in the validation set are used to calibrate a prediction interval so that its empirical coverage probability is close to its nominal coverage probability. With the estimated $\theta_{h,\alpha}$, the $100(1-\alpha)\%$ prediction interval of $d_{n+h,x}$ can be obtained as
@@
where $v$ denotes the number of observations in the validation set. As forecast horizon $h$ increases, the value of $\theta_{h,\alpha}$ generally increases.
@@ 4.5 Interval forecast evaluation metrics
For each year in the forecasting period, the $h$-step-ahead prediction intervals were computed at the $(1-\alpha)$ nominal coverage probability, where $\alpha$ denotes a level of significance. For a chosen $\alpha$, let us denote $\widehat{d}_{n+\xi|n,x}^{\text{lb}}$ and $\widehat{d}_{n+\xi|n,x}^{\text{ub}}$, as the lower and upper bounds, respectively. We compute the empirical coverage probability, which can be defined as the number of holdout data points between the lower and upper bounds for a given level of significance. It can be expressed as
@@
where the denominator is the number of curves in the forecasting period.
From the ECP, we evaluate the coverage probability difference (CPD), which is the absolute difference between empirical and nominal coverage probabilities, and use this measure to evaluate the accuracy of the interval forecast. As an absolute measure, the CPD cannot inform us about the undercoverage or overcoverage. However, it eliminates the possibility of the canceling effect, where undercoverage and overcoverage may concurrently occur. The CPD can be expressed as
@@
The smaller the value of CPD is, the more accurate the prediction interval one method provides.
The ECP and CPD measure the accuracy of prediction intervals, and neither considers the sharpness of the prediction intervals, i.e., the distance between lower and upper bounds. We also consider the interval score of Gneiting & Raftery (2007), defined as
@@
The interval score rewards a narrow prediction interval if and only if the holdout observation lies within the prediction interval. The optimal interval score is achieved when $d_{n+\xi,x}$ not only lies between the $\widehat{d}_{n+\xi|n,x}^{\text{lb}}$ and $\widehat{d}_{n+\xi|n,x}^{\text{ub}}$, but also the distance between $\widehat{d}_{n+\xi|n,x}^{\text{lb}}$ and $\widehat{d}_{n+\xi|n,x}^{\text{ub}}$ is minimal for a given age $x$.
@@ 4.6 Comparison of interval forecast accuracy
Averaging over the interval forecasts from the 24 countries, Table LABEL:tab:2 presents horizon-specific forecast errors for comparing the accuracy of the 80% prediction intervals constructed by the two transformations. Differing from the point forecast accuracy, the CDF transformation does not consistently outperform the CLR transformation, but it generally provides superior accuracy with smaller CPD and mean interval score. Further, we compare two ways of selecting the number of components in the principal component analysis for modeling the unconstrained data in the transformed space. Due to the limited space, we do not report the results for the 95% nominal coverage probability, but these can be obtained upon request from the corresponding author.
Table 2: Comparison of the interval forecast accuracy, as measured by the CPDα and mean interval score $\overline{S}_{\alpha}$, between the CLR and CDF transformations at the level of significance $\alpha=0.2$. For each transformation, we further compare two ways of selecting the number of components.
@@
Mean 1036 633 575 523 1629 930 813 648
While Table LABEL:tab:2 compares the interval forecast accuracy based on the CPD and mean interval scores, aggregated over 24 countries. Based on the averaged CPDα and $\overline{S}_{\alpha}$ across various forecast horizons, the CDF transformation provides the smaller interval forecast errors than those obtained from the CLR transformation. In Figure 5, we provide several heatmaps that show the number of times one method produces the smallest errors. For modeling the female data, the difference between the CLR and CDF transformations is marginal. Across forecast horizons, the inconsistency results between the two transformation may be because of the optimal estimation of $\theta_{h,\alpha}$ parameters in both transformations. For modeling the male data, the CDF transformation with $K=6$ is advocated with the smallest CPD and mean interval score.
@@
Figure 5: Heatmaps showing the frequency of the most accurate transformation, as measured by the CPD and mean interval score, at the nominal coverage probability of 80%.
5 Comparisons of life expectancy at birth and annuity price
An immediate application of life-table death count forecasts for individuals aged 60 and over is in the superannuation industry, whose profitability and solvency crucially rely on accurate mortality forecasts to hedge longevity risks appropriately. When a person retires, an optimal way of guaranteeing one’s financial income in retirement is to purchase an annuity product. Such an annuity is a financial contract offered by insurers guaranteeing a steady stream of payments for either a temporary or the lifetime of the annuitant in exchange for an initial premium fee.
Among many annuity products, we study temporary annuities, which have grown in popularity in a number of countries, including Australia and the USA, because lifetime immediate annuities, where rates are locked in for life, have been demonstrated to deliver poor value for money (see, e.g., Cannon & Tonks, 2008, Chapter 6). The temporary annuities pay a pre-determined and guaranteed income level higher than the income provided by a lifetime annuity for a similar amount of premium. Fixed-term annuities also offer an attractive alternative to lifetime annuities and allow the purchaser to buy a deferred annuity later.
Since the CDF transformation generally provides smaller forecast errors than those obtained from the CLR transformation. Using the CDF transformation, we obtain 50-years-ahead forecasts of life-table death counts for all countries. As a vehicle of demonstration, we display the forecasts for Australia in Figure 6 and observe a continuing trend of further longevity. These forecasts are time-series extrapolations of the past trend, and thus should be used with caution. Any extreme shocks, such as COVID-19, are likely to have a substantial impact on the mortality trend.
@@ 5 Comparisons of life expectancy at birth and annuity price
Via a standard life-table calculation, we convert the forecasts of life-table death counts $(d_{x})$ (from 2022 to 2071) to their corresponding period life expectancies ($e_{x}$). In Figure 7, we display the forecasted life expectancy at birth of the female and male populations in the 24 countries. In 2022, Japan had the highest life expectancy for females, but by 2050, the highest life expectancy may occur in Switzerland. For the male population, Switzerland had the highest life expectancy in 2022, but by 2050, the highest life expectancy may occur in Australia. Among all the countries, Bulgaria has the lowest life expectancy at birth.
@@
In Figure 8, we display the forecasted life expectancy at age 60 of the female and male populations in the 24 countries. In 2022, Japan had the highest life expectancy for females, but by 2050, the highest life expectancy may occur in Switzerland. For the male population, Australia had the highest life expectancy in 2022, and may still maintain its lead by 2050. Among all the countries, Bulgaria has the lowest life expectancy at age 60.
@@
Via a standard life-table calculation, we convert the forecasts of life-table death counts $(\bm{d}_{x})$ to their corresponding survival probabilities ($\bm{p}_{x}$) from 2022 to 2071. We plug the age-specific survival probabilities into the calculation of single-premium term immediate annuities, and we adopt a cohort approach to calculating the survival probabilities, as is the practice of insurance companies, that is, survival probability $p_{60}$ is computed from the forecasted life-table death counts in 2022, $p_{61}$ is computed from the forecasted life-table death counts in 2023, $\dots$, $p_{109}$ is computed from the forecasted life-table death counts in 2071. The $\tau$ year survival probability of a person with entry age $x$ at the contract time is determined by
$_{\tau}p_{x}=\prod^{\tau}_{j=1}p_{x+j-1}=\prod^{\tau}_{j=1}(1-q_{x+j-1})=\prod^{\tau}_{j=1}(1-\frac{d_{x+j-1}}{l_{x+j-1}}),$

### v3

@@ 5.1 Expanding window
We implement an expanding-window scheme, which is commonly used to evaluate model and parameter stability, as well as prediction accuracy over time. The expanding-window analysis assesses the stability of a model’s parameter estimates and forecasts over an expanding window across the sample. Since each data set has a different sample period, we keep the last 20 years of data for evaluation and the remaining data for training. For the 20 years of data in the testing sample, we consider one-step-ahead to 20-step-ahead forecasts to assess the model’s performance at short to medium horizons. Through the expanding-window scheme, we produce 20 one-step-ahead forecasts, 19 two-step-ahead forecasts, $\dots$, and one 20-step-ahead forecast. We compare these forecasts with the holdout data in the testing sample to determine the accuracy of the out-of-sample forecast. In Figure 6, we present a concept diagram.
@@ 5.2 Point forecast evaluation metrics
Since the age distribution of death counts can be considered a probability density function, we consider two density evaluation metrics: the symmetric discrete Kullback-Leibler divergence (KLD) (Kullback and Leibler, 1951) and the Jensen-Shannon divergence (JSD) (Shannon, 1948).
The KLD measures information loss by approximating an unknown density with its approximation. For two probability density functions, denoted by $d_{n+\xi}(u)$ and $\widehat{d}_{n+\xi|n}(u)$, the symmetric discrete KLD is defined as
@@
$=\$$\frac{1}{111\times(21-h)}\sum^{20}_{\xi=h}\sum^{111}_{x=1}[d_{n+\xi,x}(\ln d_{n+\xi,x}-\ln\widehat{d}_{n+\xi|n,x})+\widehat{d}_{n+\xi|n,x}(\ln\widehat{d}_{n+\xi|n,x}-\ln d_{n+\xi,x})],$
where $\xi$ is the forecasting period. A feature of the symmetric discrete KLD is its non-negativity.
Even though we have a sample of observed densities in the forecasting period, they may not be the actual densities. An alternative is the JSD, which can be viewed as a symmetric and smoothed version of the KLD. The JSD is defined by
@@
where $\delta_{n+\xi,x}$ measures a common quantity between $d_{n+\xi,x}$ and $\widehat{d}_{n+\xi,x}$. An example of $\delta_{n+\xi,x}$ can be its geometric mean $\delta_{n+\xi,x}=\sqrt{d_{n+\xi,x}\widehat{d}_{n+\xi,x}}$.
@@ 5.3 Comparison of point forecast accuracy
In Figure 7, we demonstrate the forecasts of the life-table death counts obtained from the CLR and CDF transformations. For one-step-ahead forecasts, the differences between the two transformations are marginal. As the forecast horizon increases to $h=20$, the difference is more visible. As measured by KLD and JSD, the CDF transformation produces smaller forecast errors than those of the CLR transformation for the $h=1$ and $h=20$ considered.
Figure 7: One-step-ahead and 20-step-ahead forecasts of life-table death counts for Australian women (first row) and men (second row). As measured by the KLD and JSD, we compute these errors obtained from the CLR and CDF transformations.
In Table LABEL:tab:KLD_tab, we compare the KLD across the 24 countries, averaged over 20 forecast horizons. Countries such as Iceland and Ireland are challenging to predict accurately using the CLR transformation. In contrast, the CDF transformation shows moderate errors. Between the two ways of selecting the number of components, it is advantageous to use $K=6$ for forecasting.
Table 2: Averaged over 20 forecast horizons, a comparison of the female and male KLD errors among the 24 countries. Heterogeneity in forecast accuracy highlights differences in prediction accuracy across the two transformations using our chosen forecasting method. Eigenvalue ratio (EVR) is a criterion for selecting the number of components.
Female Male
EVR $K=6$ EVR $K=6$
Country CLR CDF CLR CDF CLR CDF CLR CDF
AUS 0.0117 0.0091 0.0040 0.0029 0.0208 0.0195 0.0096 0.0047
AUT 0.0092 0.0056 0.0101 0.0053 0.0186 0.0074 0.0334 0.0071
BEL 0.0144 0.0059 0.0117 0.0051 0.0509 0.0305 0.0209 0.0085
BGR 0.0771 0.0335 0.0220 0.0242 0.0527 0.0220 0.0190 0.0159
CAN 0.0066 0.0052 0.0051 0.0035 0.0355 0.0340 0.0147 0.0102
CZE 0.0406 0.0123 0.0106 0.0137 0.1508 0.0244 0.0115 0.0142
DEN 0.0176 0.0052 0.0142 0.0067 0.0380 0.0179 0.0285 0.0184
FIN 0.0481 0.0177 0.0191 0.0053 0.0991 0.0468 0.0351 0.0109
FRA 0.0250 0.0081 0.0162 0.0074 0.0460 0.0333 0.0242 0.0247
HUN 0.0750 0.0060 0.0074 0.0068 0.1154 0.0618 0.0655 0.0434
ICE 0.3189 0.0696 0.3708 0.0567 0.1629 0.0465 0.2282 0.0475
IRE 0.1731 0.0125 0.0121 0.0090 0.2546 0.0650 0.0455 0.0230
ITA 0.0104 0.0223 0.0067 0.0143 0.0442 0.0280 0.0201 0.0205
JPN 0.0954 0.0846 0.1072 0.0805 0.0075 0.0129 0.0096 0.0120
NLD 0.0303 0.0134 0.0134 0.0101 0.0602 0.0334 0.0392 0.0306
NOR 0.0323 0.0160 0.0170 0.0068 0.0798 0.0439 0.0362 0.0208
NZ 0.0209 0.0075 0.0091 0.0057 0.2086 0.0290 0.0176 0.0095
PRT 0.0090 0.0063 0.0076 0.0053 0.0180 0.0095 0.0284 0.0091
SLO 0.1185 0.0257 0.0214 0.0234 0.0794 0.0537 0.0430 0.0297
SPA 0.0217 0.0171 0.0124 0.0089 0.0304 0.0112 0.0302 0.0058
SWE 0.0295 0.0101 0.0154 0.0166 0.0351 0.0040 0.0268 0.0047
SWI 0.0083 0.0068 0.0062 0.0077 0.0190 0.0108 0.0069 0.0058
UK 0.0175 0.0068 0.0100 0.0045 0.0342 0.0293 0.0082 0.0055
USA 0.0086 0.0073 0.0082 0.0061 0.0172 0.0170 0.0158 0.0192
Mean 0.0508 0.0173 0.0308
0.0140
0.0700 0.0288 0.0341
0.0167
Averaging over point forecast errors obtained from the 24 countries, Figure 8 presents horizon-specific plots to compare the accuracy of the two transformations, as measured by the KLD and JSD. Between the CLR and CDF transformations, the CDF transformation consistently outperforms the CLR transformation. This superiority is credited to the monotonicity constraint enjoyed by the CDF transformation. For modeling the unconstrained data in Euclidean space, it is advantageous to use $K=6$ for forecasting. In contrast, the number of components selected by the EVR tends to be fewer than optimal. Choosing a larger number of components results in only a minor loss, whereas selecting too few components leads to a more substantial loss.
Figure 8: Comparison of the point forecast accuracy, measured by the KLD and JSD, between the CLR and CDF transformations. For each transformation, we perform a principal component analysis, selecting the number of components using the EVR criterion or setting $K=6$.
@@ 5.4 Construction of pointwise prediction intervals
Using the data in the training sample, we applied an expanding-window approach to obtain $h$-step-ahead density forecasts in the validation set for $h=1,\dots,19$. For different forecast horizons, we have different numbers of curves in the validation set. For $h=1$, we have 20 years to evaluate forecast errors; for $h=19$, we have two years to evaluate the residual functions between the holdout data in the validation set and their forecasts. Based on these residuals, we can compute the functional standard deviation, which requires at least two years of data. The forecast errors can be denoted by $\widehat{\epsilon}_{\nu,x}=d_{\nu,x}-\widehat{d}_{\nu,x}$, where $\widehat{d}_{\nu,x}$ denotes the forecasts obtained from the two transformations, and $\nu=1,\dots,(21-h)$ denotes the number of observations in the validation set for $h=1,\dots,19$. When $h=1$, we have 20 observations in the validation set; when $h=19$, we have two observations.
Let us denote $\Gamma_{\nu,x}=\text{sd}(\widehat{\epsilon}_{\nu,x})$, where $\text{sd}(\cdot)$ can be pointwise standard deviation. We denote $\bm{\Gamma}_{\nu}=(\Gamma_{\nu,1},\dots,\Gamma_{\nu,111})^{\top}$ and $\bm{\widehat{\epsilon}}_{\nu}=(\widehat{\epsilon}_{\nu,1},\dots,\widehat{\epsilon}_{\nu,111})^{\top}$. For a level of significance $\alpha$, commonly $\alpha=0.2$ or 0.05, our aim is to find $\theta_{h,\alpha}$ such that approximately $100(1-\alpha)\%$ of the residuals satisfy
@@
According to the law of large numbers, one can achieve the following
@@
To determine the optimal $\theta_{h,\alpha}$, the samples in the validation set are used to calibrate a prediction interval so that its empirical coverage probability is close to its nominal coverage probability (see also Shang and Haberman, 2026). With the estimated $\theta_{h,\alpha}$, the $100(1-\alpha)\%$ prediction interval of $d_{n+h,x}$ can be obtained as
@@
where $v$ denotes the number of observations in the validation set. As the forecast horizon $h$ increases, the value of $\theta_{h,\alpha}$ generally increases, reflecting an increasing degree of forecast uncertainty.
@@ 5.5 Interval forecast evaluation metrics
For each year in the forecasting period, the $h$-step-ahead prediction intervals were computed at the $(1-\alpha)$ nominal coverage probability, where $\alpha$ denotes a level of significance. For a chosen $\alpha$, let us denote $\widehat{d}_{n+\xi|n,x}^{\text{lb}}$ and $\widehat{d}_{n+\xi|n,x}^{\text{ub}}$, as the lower and upper bounds, respectively. We compute the empirical coverage probability, defined as the proportion of holdout data points that fall within the lower and upper bounds for a given significance level. It can be expressed as
@@
where the denominator is the number of curves in the forecasting period, which depends on the forecast horizon.
From the ECP, we compute the coverage probability difference (CPD), the absolute difference between the empirical and nominal coverage probabilities, and use it to assess the accuracy of the interval forecast. As an absolute measure, the CPD cannot inform us about undercoverage or overcoverage. However, it eliminates the possibility of the canceling effect, in which undercoverage and overcoverage may occur concurrently. The CPD can be expressed as
@@
The smaller the value of CPD, the more accurate the prediction interval provided by one method.
The ECP and CPD measure the accuracy of prediction intervals, and neither considers the sharpness of the prediction intervals, i.e., the distance between lower and upper bounds. We also consider the interval score of Gneiting and Raftery (2007), defined as
@@
The interval score rewards a narrow prediction interval if and only if the holdout observation lies within the prediction interval. The optimal interval score is achieved when $d_{n+\xi,x}$ is not only between $\widehat{d}_{n+\xi|n,x}^{\text{lb}}$ and $\widehat{d}_{n+\xi|n,x}^{\text{ub}}$, but the distance between $\widehat{d}_{n+\xi|n,x}^{\text{lb}}$ and $\widehat{d}_{n+\xi|n,x}^{\text{ub}}$ is minimal for a given age $x$.
@@ 5.6 Comparison of interval forecast accuracy
Averaging over the interval forecasts from the 24 countries, Table LABEL:tab:2 presents horizon-specific forecast errors for comparing the accuracy of the 80% prediction intervals constructed by the two transformations. Differing from point forecast accuracy, the CDF transformation does not consistently outperform the CLR transformation; however, it generally provides superior accuracy with smaller CPD and mean interval scores. Furthermore, we compare two methods for selecting the number of components in principal component analysis for modeling unconstrained data in the transformed space. Due to limited space, we do not report the results for the 95% nominal coverage probability; however, these can be obtained upon request from the corresponding author.
Table 3: Comparison of the interval forecast accuracy, as measured by the CPDα and mean interval score $\overline{S}_{\alpha}$, between the CLR and CDF transformations at the level of significance $\alpha=0.2$. For each transformation, we compare two methods for selecting the number of components.
@@
Mean 1036 633 575
523
1629 930 813
648
@@
Figure 9: Heatmaps showing the frequency of the most accurate transformation, as measured by the CPD and mean interval score, at the nominal coverage probability of 80%. Each row sums up to 24 countries.
In Table LABEL:tab:2, we compare the interval forecast accuracy using the CPD and mean interval scores, aggregated across 24 countries. Based on the averaged CPDα and $\overline{S}_{\alpha}$ across various forecast horizons, the CDF transformation generally provides smaller interval forecast errors than those obtained from the CLR transformation. In Figure 9, we present several heatmaps illustrating the frequency with which each method yields the smallest errors. For modeling the female data, the difference between the CLR and CDF transformations is marginal. Across forecast horizons, the inconsistency between the two transformations may stem from the optimal estimation of the parameters $\theta_{h,\alpha}$ in both. For modeling male data, the CDF transformation with $K=6$ is recommended, as it yields the smallest CPD and the smallest mean interval score.
6 Multi-country comparisons
6.1 Comparison of life expectancy at birth
As a widely used summary measure for both demographers and actuaries, we evaluate and compare forecast life expectancies at birth across countries (see also Shang et al., 2011). Using the LifeTable function of the MortalityLaws package (Pascariu, 2025) in [Uncaptioned image] , we convert the forecast age distribution of deaths into life expectancy at birth.
@@ 6.1 Comparison of life expectancy at birth
The CDF transformation generally yields smaller forecast errors than the CLR transformation. Using the CDF transformation, we obtain 50-years-ahead forecasts of life-table death counts for all countries. As a demonstration, we display the forecasts for Australia in Figure 10 and observe a continuing trend of greater longevity. These forecasts are time-series extrapolations of past trends and should be used with caution. Any extreme shocks, such as the COVID-19 pandemic, are likely to have a substantial impact on the mortality trend.
Via a standard life-table calculation, we convert the forecasts of life-table death counts $(d_{x})$ (from 2022 to 2071) to their corresponding period life expectancies ($e_{x}$). In Figure 11, we display the forecasted life expectancy at birth of the female and male populations in the 24 countries. In 2022, Japan had the highest life expectancy for females, but by 2050, the highest may be in Switzerland. In 2022, Switzerland had the highest life expectancy among men, but by 2050, Australia may have the highest. Among all countries, Bulgaria has the lowest life expectancy at birth, due to factors such as public health, socioeconomic status, and healthcare.
@@
In Figure 12, we display the forecasted life expectancy at age 60 of the female and male populations in the 24 countries. In 2022, Japan had the highest life expectancy for females, but by 2050, the highest may be in Switzerland. Among males, Australia had the highest life expectancy in 2022 and may still maintain its lead by 2050. Among all countries, Bulgaria has the lowest life expectancy at 60 years of age.
@@
6.2 Comparison of annuity price
An immediate application of life-table death count forecasts for individuals aged 60 and over is in the superannuation industry, whose profitability and solvency crucially depend on accurate mortality forecasts to appropriately hedge longevity risks. When a person retires, a reliable way to ensure a steady financial income is to purchase an annuity. Such an annuity is a financial contract offered by insurers that guarantees a steady stream of payments for either a temporary or the lifetime of the annuitant in exchange for an initial premium.
Among many annuity products, we study temporary annuities, which have grown in popularity in a number of countries, including Australia and the USA, because lifetime immediate annuities, where rates are locked in for life, have been demonstrated to deliver poor value for money (see, e.g., Cannon and Tonks, 2008, Chapter 6). Temporary annuities pay a predetermined, guaranteed income level higher than that provided by a lifetime annuity for a similar premium amount. Fixed-term annuities also offer an attractive alternative to lifetime annuities, allowing the purchaser to purchase a deferred annuity later.
Via a standard life-table calculation, we convert the forecasts of life-table death counts $(\bm{d}_{x})$ to their corresponding survival probabilities ($\bm{p}_{x}$) from 2022 to 2071. We plug the age-specific survival probabilities into the calculation of single-premium term immediate annuities, and we adopt a cohort approach to calculating the survival probabilities, as is the practice of insurance companies, that is, survival probability $p_{60}$ is computed from the forecast life-table death counts in 2022, $p_{61}$ is computed from the forecast life-table death counts in 2023, $\dots$, $p_{109}$ is computed from the forecast life-table death counts in 2071. The $\tau$ year survival probability of a person with entry age $x$ at the contract time is determined by
$_{\tau}p_{x}=\prod^{\tau}_{j=1}p_{x+j-1}$ $=\prod^{\tau}_{j=1}(1-q_{x+j-1})$
$=\prod^{\tau}_{j=1}(1-\frac{d_{x+j-1}}{l_{x+j-1}}),$